# ORCID ID: ]https://orcid.org/$0$-0002-9154-1878

### v1

@@ Abstract
The model of localized fermions on the triangular lattice is analyzed in means of the Monte Carlo simulations in the grand canonical ensemble. The Hamiltonian of the system has a form of the extended Hubbard model (at the atomic limit) with nearest-neighbor Ising-like magnetic $J$ interactions and onsite Coulomb $U$ interactions. The model is investigated for both signs of $J$, arbitrary $U$ interaction and arbitrary chemical potential $\mu$ (or, equivalently, arbitrary particle concentration $n$). Based on the specific heat capacity and sublattice magnetization analyses, the phase diagrams of the model are determined. For ferromagnetic case ($J<0$), the transition from the ordered phase (which is a standard ferromagnet and can be stable up to $k_{B}T/|J|\approx 0.61$) is found to be second-order (for sufficiently large temperatures $k_{B}T/|J|\gtrsim 0.2$) or first-order (for $-1<U/|J|<-0.65$ at the half-filling, i.e., $n=1$). In the case of $J>0$, the ordered phase occurs in a range of $-1/2<U/|J|<0$ (for $n=1$), while for larger $U$ the state with short-range order is also found (also for $n\neq 1$). The ordered phase is characterized by an antiferromagnetic arrangement of magnetic moments in two sublattices forming the hexagonal lattice. The transition from this ordered phase, which is found also for $\mu\neq 0$ ($n\neq 1$) and $U/|J|>-1/2$ is always second-order for any model parameters. The ordered phase for $J>0$ can be stable up to $k_{B}T/|J|\approx 0.06$.
@@ I Introduction
The Ising model $\hat{H}_{I}=\sum_{i,j}J_{i,j}\hat{s}_{i}\hat{s}_{j}$ is one of the simplest model which can describe the magnetism in the system of localized magnetic moments. Its exact solutions with interactions restricted to the nearest-neighbors (i.e., $J_{ij}\equiv J$ if $i$ and $j$ are neighboring site and $J_{ij}\equiv 0$ otherwise) for the one-dimensional chain as well as the two-dimensional square lattice for both signs of the interaction are very well-known [7, 8, 9]. The model on the triangular lattice has been also extensively studied. As that lattice has a geometrical frustration, the system exhibits different behaviors, which are dependent on the sign of interaction $J$. Rigorous results for this lattice [10, 11] show that the transition temperature is $k_{B}T_{c}/|J|=0.607$ for ferromagnetic case ($J<0$), whereas for antiferromagnetic case ($J>0$) the transition is absent and there is no long-range order in the system at any temperature. Within the use of Bethe-Peierls approximation [12] and Monte Carlo simulations [13], it was shown that ordered phase (with so-called $\sqrt{3}\times\sqrt{3}$ order) can occur in the model with $J>0$ for non-zero fields or for large enough next-nearest-neighbor interactions in finite temperatures (cf. also mean-field solutions [14]). With longer-range interactions even more complex ordered patterns are possible [15, 16, 17].
@@
where $\hat{c}_{i\sigma}^{{\dagger}}$ ($\hat{c}_{i\sigma}^{\ }$) denotes the creation (annihilation) operator of a fermion with spin $\sigma$ ($\sigma\in\{\uparrow,\downarrow\}$) at lattice site $i$ and other operators are defined as $\hat{s}_{i}=\left(\hat{n}_{i\uparrow}-\hat{n}_{i\downarrow}\right)/2$, $\hat{n}_{i}=\hat{n}_{i\uparrow}+\hat{n}_{i\downarrow}$, $\hat{n}_{i\sigma}=\hat{c}_{i\sigma}^{{\dagger}}\hat{c}_{i\sigma}^{\ }$. $\sum_{\langle i,j\rangle}$ is a summation over nearest-neighbor sites independently and $z$ is a coordination number (number of the nearest neighbors). $U$ denotes the onsite Hubbard interaction and $J$ is the Ising-like magnetic interaction between nearest-neighbor sites $i$ and $j$. Finally, $\mu$ stands for the chemical potential (shifted by $-U/2$), which controls total particle concentration $n$ in the system (for $\mu=0$ one gets the half-filling $n=1$[31, 32].
@@
Model (1) was extensively studied for hyper-cubic lattices. In such a case there is a full correspondence between $J<0$ and $J>0$ cases. In the variational approach with the mean-field approximation for the intersite term and rigorous treatment of the single-site term [33, 34, 35, 32] (rigorous solution for $z\rightarrow+\infty$), the phase transition between magnetically ordered phases (ferromagnetic or antiferromagnetic depending on the sign of $J$ interaction) was found to be first-order (discontinuous) for $k_{B}T/|J|<1/3$ and second-order (continuous) for $k_{B}T/|J|>1/3$. The tricritical point, where transition changes its order at half-filling is located at $U/|J|=(2/3)\ln 2$ [34, 32]. The Monte Carlo simulations were performed for the model on two-dimensional square lattice with $z=4$ [36, 37, 38]. It was found that the structure of the phase diagrams is similar as in the variational approach with reduction of the range of ordered phase occurrence, e.g., the transition temperatures are reduced if compared with $z\rightarrow+\infty$ limit. Also some exact results for one-dimensional chain ($z=2$) were obtained [39, 40].
Figure 1: Phase diagrams for the model with ferromagnetic ($J=-1$) interactions, shown for three different lattice sizes ($L=12^{2},24^{2},36^{2}$). Triangle markers indicate first-order phase transitions, while circular markers represent second-order phase transitions. In the bottom-left corner, Monte Carlo averaged spin configurations for the identified ferromagnetic (F) phase at $U=-0.6$, $\mu=0.0$, and $k_{B}T=0.15$ and non-ordered (NO) phase at $U=-0.6$, $\mu=0.0$, and $k_{B}T=0.4$, are shown for a lattice of size $L=6\times 6$.
@@ II Method description: Monte Carlo algorithm
Figure 2: Specific heat capacities (top row), sublattice magnetizations (second row), and the energy distribution functions near the phase transition on a $24\times 24$ lattice (third line) for ferromagnetic interaction ($J=-1$) and representative values of $U$ ($U=-1.0,-0.7,-0.6,5.0$, from the left) at $\mu=0$.
@@ III Results and discussion
In this section, the results of the simulations for both signs of the intersite magnetic interaction are presented. We present the phase diagrams, which are determined by careful examination of the temperature dependence of specific heat capacity $C=dE/dT$ ($E$ the energy of the system per lattice site) and sublattice magnetization $m_{\alpha}$. In particular, the specific heat capacity is determined using the formula $C=\langle\varepsilon^{2}\rangle-\langle\varepsilon\rangle^{2}/T^{2}$, where $\langle\varepsilon\rangle$ ($\langle\varepsilon^{2}\rangle$) represents the averages of energies $\varepsilon$ ($\varepsilon^{2}$) sampled by the Monte Carlo algorithm. Moreover, magnetization $m_{\alpha}$ is an average magnetic moment (per site, multiplied by the factor of $2$) in $\alpha$ sublattice ($\alpha\in\{A,B,C\}$), which build the triangular lattice, cf., e.g., [12, 2, 29, 30, 31] ($m_{\alpha}=(6/L)\sum_{i\in\alpha}\langle\hat{s}_{i}\rangle$).
Figure 3: Specific heat capacities (top row), sublattice magnetizations (second row), number of particles (third row), and the energy distribution functions near the phase transition on a $24\times 24$ lattice (fourth line) for ferromagnetic interaction ($J=-1$) and representative values of $U$ and $\mu$ (from the left: $U=-0.7$ and $\mu=-0.15,-0.10$; $U=-0.6$ and $\mu=-0.25,-0.18,-0.12$).
In the present work, all energies (i.e., $k_{B}T$, $U$, and $\mu$) are shown in units of $|J|=z\times(|J|/z)$ (the notation of Eq. (1)) in the full correspondence to the results previously presented in Refs. [33, 34, 35, 32, 36, 37, 38, 31]. Thus, in principle, we use $|J|=1$. Because the model exhibits the particle-hole symmetry [32, 31], $k_{B}T$-$\mu$ phase diagrams are symmetric with respect to $\mu=0$ and they are presented only for $\mu\leq 0$ (or, equivalently, for $n\leq 1$).
@@ III.1 Ferromagnetic interactions
First, we discuss the case of ferromagnetic intersite interactions (i.e., $J<0$). In this case, there is no frustration in the system, and we have identified two phases occurring in it: (i) the non-ordered (NO) phase which is defined by zero average magnetization in each sublattice ($m=m_{\alpha}=0$, i.e., equal numbers of particles with spin-$\uparrow$ and with spin-$\downarrow$ in any of the sublattices) and (ii) the ferromagnetic (F) phase in which $m=m_{\alpha}$ and $|m|>0$ (all three sublattices have the same magnetization). In both phases, concentrations of the particles in each sublattice are the same, i.e., $n=n_{\alpha}$. Note that, in this case, the spin configurations with $m=|m|$ (all averaged moments $\uparrow$-direction) and $m=-|m|$ (all all averaged moments $\downarrow$-direction) are equivalent (because there is no external magnetic field in the system).
@@
Let us begin our discussion with the study of the influence of the Hubbard interaction $U$ at zero chemical potential $\mu=0$ (which corresponds to the half-filling in model (1), $n=1$), for which the phase diagrams are depicted in Fig. 1 (the left top panel). One can see that the ferromagnetic interaction $J$ leads to a relatively simple $k_{B}T$$U$ phase diagram with a phase transition from the non-ordered phase (NO) to the ordered (F) phase for $U>-1$, moving towards higher temperatures $k_{B}T$ with increasing $U$. At sufficiently large on-site density interaction $U$ ($U>3$), it reaches an almost constant value of $k_{B}T\approx 0.605$, which aligns with the exact solution of the classical Ising model, to which our model can be mapped in the limit of large $U$ (for $\mu=0$).
@@
These phase transitions are obtained from the analysis of specific heat capacity and sublattice magnetizations on finite lattices with $L=12\times 12$ to $L=36\times 36$ sites, which are illustrated in Fig. 2 for representative values of $U$ ($U=-1.0,-0.7,-0.6,5.0$). One can see that in the case of $U\leq-1$ (the NO phase), we have found a broad maximum in the specific heat capacity curves which does not change with increasing $L$, and there is also no change in sublattice magnetizations, which equal zero. This can be attributed to the presence of a Schottky anomaly, which indicates a small number of discrete energy levels that dominate the system behavior and quantify the spacing between these levels [46]. For $U>-1$ (where the F phase is found at low temperatures), we have found the typical $\lambda$-like singularity behavior of specific heat capacity, which increases its maximum value with the increasing $L$ as well as the rapid change in sublattice magnetizations from $|m|=0$ to $|m|=1$ with the decreasing temperature. To distinguish the order of these phase transitions, we have used the method by Challa et al. [47, 48, 49, 50], which is based on the energy distribution functions $P(E)$ for temperatures near phase transition. For $U$ smaller than $U\approx-0.65$ (temperatures smaller than $k_{B}T\approx 0.2$), the energy distribution function exhibits a two-peaked structure near the phase transition (as shown in Fig. 2 for $U=-0.7$ and $k_{B}T\approx 0.19$), indicating a first-order phase transition from the NO phase to the F phase with decreasing temperature. This double-peak distribution indicates that states with two different energies of the system are equally probable at the transition temperature and one of them corresponds to the metastable solution in the vicinity of the transition line (however, the double-peak structure resembles a wide peak for model parameters shown). On the other hand, for higher $U$, only one peak in $P(E)$ is observed, which denotes a second-order phase transition (absence of metastability in the system) as shown for $U=-0.6$ and $k_{B}T\approx 0.28$. We also presents the results for $U=5.0$, where we get the second-order transition at $k_{B}T\approx 0.604$, which is in very good accordance with the exact result for the Ising model ($k_{B}T\approx 0.607$) [10, 11]. However, the fluctuations are more visible due to higher temperatures and more Monte Carlo steps would provide smoother curves.
@@
Now we are ready to analyze the behavior of the system with changing $\mu$ (for $n\neq 1$). Exemplary $k_{B}T$$\mu$ phase diagrams for different values of $U$ are shown in Fig. 1. For all values of $U$ one sees the universal behavior of F–NO transition temperature, namely, the increase of $|\mu|$ leads to a decrease of the temperature. In the case of $U=-0.7$, the transition remains of first-order across the entire range of the parameter $\mu$. Another behavior is observed for $U$ values larger than $U\approx-0.65$, where a second-order transition occurs near $\mu=0$ and, with increasing $|\mu|$, changes its order into a first-order transition (cf. diagrams for $U=-0.6$ and $U=-0.5$). With further increasing of onsite repulsion $U$, the region of the F phase occurrence extends and the first-order line becomes almost vertical (cf. $U=0.0$ case). It is remarkable that the temperature at which the F-NO transition changes its order (the tricritical point) is almost independent of $U$ and $\mu$, remaining around $k_{B}T\approx 0.18-0.20$, which is in an agreement with previous studies [32, 31].
In Fig. 3, we present dependence of several quantities with increasing temperature away half-filling to give an insight and justify the phase diagrams discussed above. The first and third columns ($U=-0.7$, $\mu=-0.15$ and $U-0.6$, $\mu=-0.25$) present the dependencies in the NO phase with the broad peak in the specific heat capacity accompanied by zero sublattice magnetizations and continuous decrease in number of particles with decreasing temperature. The second and the fourth columns ($U-0.7$, $\mu=-0.10$ and $U=-0.6$, $\mu=-0.18$) clearly indicate that, with increasing temperature, the first-order F-NO transition occurs, with a sharp (discontinuous) change of $m$ and $n$ at the transition temperature, and double peak-structure in the energy distribution $P(E)$ near the transition (which is more clearly visible than previously discussed example in Fig. 2). Finally, it is clearly seen that for $U=-0.6$ and $\mu=-0.12$ (the fifth column) the F-NO transition is second-order. Note that the particle concentration $n$ approaches zero in the NO phase at the ground state, whereas in the F phase it goes to $1$ as it is expected for the exact ground-state results [32, 31].
It is worth to mention that the findings for $J=-1$ are in a very good agreement, at least qualitatively, with the results obtained by the mean-field approach presented in [32, 31]. Obviously, the mean-field overestimates the transition temperatures. Moreover, the results for the ferromagnetic coupling are also very similar (qualitatively and quantitatively) to those obtained for the square lattice [36, 37, 38, 32]. The predicted transition temperatures for the triangular lattice ($z=6$) are slightly higher than those for the square lattice ($z=4$), which is due to the higher number of the nearest-neighbors in the former case.

### v2

# ORCID ID: ]https://orcid.org/0000-0002-9154-1878

@@ Abstract
The model of localized fermions on the triangular lattice is analyzed in means of the Monte Carlo simulations in the grand canonical ensemble. The Hamiltonian of the system has a form of the extended Hubbard model (at the atomic limit) with nearest-neighbor Ising-like magnetic $J$ interactions and onsite Coulomb $U$ interactions. The model is investigated for both signs of $J$, arbitrary $U$ interaction and arbitrary chemical potential $\mu$ (or, equivalently, arbitrary particle concentration $n$). Based on the specific heat capacity and sublattice magnetization analyses, the phase diagrams of the model are determined. For ferromagnetic case ($J<0$), the transition from the ordered phase (which is a standard ferromagnet and can be stable up to $k_{B}T/|J|\approx 0.61$) is found to be second-order (for sufficiently large temperatures $k_{B}T/|J|\gtrsim 0.2$) or first-order (for 1<U/|J|<0.65-1<U/|J|<-0.65 at the half-filling, i.e., $n=1$). In the case of $J>0$, the ordered phase occurs in a range of 1/2<U/|J|<0-1/2<U/|J|<0 (for $n=1$), while for larger $U$ the state with short-range order is also found (also for $n\neq 1$). The ordered phase is characterized by an antiferromagnetic arrangement of magnetic moments in two sublattices forming the hexagonal lattice. The transition from this ordered phase, which is found also for $\mu\neq 0$ ($n\neq 1$) and U/|J|>1/2U/|J|>-1/2 is always second-order for any model parameters. The ordered phase for $J>0$ can be stable up to $k_{B}T/|J|\approx 0.06$.
@@ I Introduction
The Ising model $\hat{H}_{I}=\sum_{i,j}J_{i,j}\hat{s}_{i}\hat{s}_{j}$ is one of the simplest model which can describe the magnetism in the system of localized magnetic moments. Its exact solutions with interactions restricted to the nearest-neighbors (i.e., $J_{ij}\equiv J/z$ if $i$ and $j$ are neighboring site and $J_{ij}\equiv 0$ otherwise) for the one-dimensional chain as well as the two-dimensional square lattice for both signs of the interaction are very well-known [7, 8, 9]. The model on the triangular lattice has been also extensively studied. As that lattice has a geometrical frustration, the system exhibits different behaviors, which are dependent on the sign of interaction $J$. Rigorous results for this lattice [10, 11] show that the transition temperature is $k_{B}T_{c}/|J|=0.607$ for ferromagnetic case ($J<0$), whereas for antiferromagnetic case ($J>0$) the transition is absent and there is no long-range order in the system at any temperature. Within the use of Bethe-Peierls approximation [12] and Monte Carlo simulations [13], it was shown that ordered phase (with so-called $\sqrt{3}\times\sqrt{3}$ order) can occur in the model with $J>0$ for non-zero fields or for large enough next-nearest-neighbor interactions in finite temperatures (cf. also mean-field solutions [14]). With longer-range interactions even more complex ordered patterns are possible [15, 16, 17].
@@
where $\hat{c}_{i\sigma}^{{\dagger}}$ ($\hat{c}_{i\sigma}^{\ }$) denotes the creation (annihilation) operator of a fermion with spin $\sigma$ ($\sigma\in\{\uparrow,\downarrow\}$) at lattice site $i$ and other operators are defined as $\hat{s}_{i}=\left(\hat{n}_{i\uparrow}-\hat{n}_{i\downarrow}\right)/2$, $\hat{n}_{i}=\hat{n}_{i\uparrow}+\hat{n}_{i\downarrow}$, $\hat{n}_{i\sigma}=\hat{c}_{i\sigma}^{{\dagger}}\hat{c}_{i\sigma}^{\ }$. $\sum_{\langle i,j\rangle}$ is a summation over nearest-neighbor sites independently and $z$ is a coordination number (number of the nearest neighbors). $U$ denotes the onsite Hubbard interaction and $J$ is the Ising-like magnetic interaction between nearest-neighbor sites $i$ and $j$. Finally, $\mu$ stands for the chemical potential (shifted by U/2-U/2), which controls total particle concentration $n$ in the system (for $\mu=0$ one gets the half-filling n=1n=1[31, 32].
@@
Figure 1: Schematic three dimensional phase diagrams (in $\mu$-$U$-$k_{B}T$ space) for ferromagnetic case ($J<0$, the left panel) and antiferromagnetic case ($J>0$, the right panel). The blue and red surfaces denote first order (discontinuous) and second order (continuous) transitions.
Model (1) was extensively studied for hyper-cubic lattices. In such a case there is a full correspondence between $J<0$ and $J>0$ cases. In the variational approach with the mean-field approximation for the intersite term and rigorous treatment of the single-site term [33, 34, 35, 32] (rigorous solution for $z\rightarrow+\infty$), the phase transition between magnetically ordered phases (ferromagnetic or antiferromagnetic depending on the sign of $J$ interaction) was found to be first-order (discontinuous) for $k_{B}T/|J|<1/3$ and second-order (continuous) for $k_{B}T/|J|>1/3$. The tricritical point, where transition changes its order at half-filling is located at $U/|J|=(2/3)\ln 2$ [34, 32]. The Monte Carlo simulations were performed for the model on two-dimensional square lattice with $z=4$ [36, 37, 38]. It was found that the structure of the phase diagrams is similar as in the variational approach with reduction of the range of ordered phase occurrence, e.g., the transition temperatures are reduced if compared with $z\rightarrow+\infty$ limit. Also some exact results for one-dimensional chain ($z=2$) were obtained [39, 40]. Unfortunately, any rigorous and/or analytical results for model (1) are not known in dimensions $1<d<+\infty$ up to now (apart from those in the mean-field approximation).
@@ II Method description: Monte Carlo algorithm
Figure 2: Phase diagrams for the model with ferromagnetic (J=1J=-1) interactions, shown for three different lattice sizes ($L=12^{2},24^{2},36^{2}$). Triangle markers indicate first-order phase transitions, while circular markers represent second-order phase transitions. In the bottom-left corner, Monte Carlo averaged spin configurations for the identified ferromagnetic (F) phase at U=0.6U=-0.6, $\mu=0.0$, and $k_{B}T=0.15$ and non-ordered (NO) phase at U=0.6U=-0.6, $\mu=0.0$, and $k_{B}T=0.4$, are shown for a lattice of size $L=6\times 6$.
Figure 3: Specific heat capacities (top row), sublattice magnetizations (second row), and the energy distribution functions near the phase transition on a $24\times 24$ lattice (third line) for ferromagnetic interaction (J=1J=-1) and representative values of $U$ (U=1.0,0.7,0.6,5.0U=-1.0,-0.7,-0.6,5.0, from the left) at $\mu=0$.
@@ III Results and discussion
In this section, the results of the simulations for both signs of the intersite magnetic interaction are presented. We present the phase diagrams, which are determined by careful examination of the temperature dependence of specific heat capacity $C=dE/dT$ ($E$ - the energy of the system per lattice site) and sublattice magnetization $m_{\alpha}$. In particular, the specific heat capacity is determined using the formula $C=\langle\varepsilon^{2}\rangle-\langle\varepsilon\rangle^{2}/T^{2}$, where $\langle\varepsilon\rangle$ ($\langle\varepsilon^{2}\rangle$) represents the averages of energies $\varepsilon$ ($\varepsilon^{2}$) sampled by the Monte Carlo algorithm. Moreover, magnetization $m_{\alpha}$ is an average magnetic moment (per site, multiplied by the factor of $2$) in $\alpha$ sublattice ($\alpha\in\{A,B,C\}$), which build the triangular lattice, cf., e.g., [12, 2, 29, 30, 31] ($m_{\alpha}=(6/L)\sum_{i\in\alpha}\langle\hat{s}_{i}\rangle$).
Figure 4: Specific heat capacities (top row), sublattice magnetizations (second row), number of particles (third row), and the energy distribution functions near the phase transition on a $24\times 24$ lattice (fourth line) for ferromagnetic interaction (J=1J=-1) and representative values of $U$ and $\mu$ (from the left: U=0.7U=-0.7 and $\mu=-0.15,-0.10$; U=0.6U=-0.6 and $\mu=-0.18,-0.12$).
In the present work, all energies (i.e., $k_{B}T$, $U$, and $\mu$) are shown in units of $|J|=z\times(|J|/z)$ (the notation of Eq. (1)) in the full correspondence to the results previously presented in Refs. [33, 34, 35, 32, 36, 37, 38, 31]. Thus, in principle, we use $|J|=1$. Because the model exhibits the particle-hole symmetry [32, 31], $k_{B}T$-$\mu$ phase diagrams are symmetric with respect to $\mu=0$ and they are presented only for $\mu\leq 0$ (or, equivalently, for $n\leq 1$). Fig. 1 shows schematic phase diagrams of model (1) for both signs for $J$ interaction and illustrates the positions of phase boundaries of ordered phases (defined in the next section) of first (red) and second (blue) order in the $\mu-U-T$ space. In the figure for $J>0$ case, the region of the occurrence of the short-range order inside the non-ordered phase is not indicated.
@@ III.1 Ferromagnetic interactions
First, we discuss the case of ferromagnetic intersite interactions (i.e., $J<0$). In this case, there is no frustration in the system, and we have identified two phases occurring in it: (i) the non-ordered (NO) phase which is defined by zero average magnetization in each sublattice ($m=m_{\alpha}=0$), indicating equal average occupancies of spin-$\uparrow$ and spin-$\downarrow$ particles and the absence of both long-range and significant short-range magnetic order and (ii) the ferromagnetic (F) phase in which $m=m_{\alpha}$ and $|m|>0$ (all three sublattices have the same magnetization). In both phases, concentrations of the particles in each sublattice are the same, i.e., $n=n_{\alpha}$. Note that, in this case, the spin configurations with $m=|m|$ (all averaged moments $\uparrow$-direction) and m=|m|m=-|m| (all all averaged moments $\downarrow$-direction) are equivalent (because there is no external magnetic field in the system).
@@
Let us begin our discussion with the study of the influence of the Hubbard interaction $U$ at zero chemical potential $\mu=0$ (which corresponds to the half-filling in model (1), $n=1$), for which the phase diagrams are depicted in Fig. 2 (the left top panel). One can see that the ferromagnetic interaction $J$ leads to a relatively simple $k_{B}T$$U$ phase diagram with a phase transition from the non-ordered phase (NO) to the ordered (F) phase for U>1U>-1, moving towards higher temperatures $k_{B}T$ with increasing $U$. At sufficiently large on-site density interaction $U$ ($U>3$), it reaches an almost constant value of $k_{B}T\approx 0.605$, which aligns with the exact solution of the classical Ising model, to which our model can be mapped in the limit of large $U$ (for $\mu=0$).
@@
These phase transitions are obtained from the analysis of specific heat capacity and sublattice magnetizations on finite lattices with $L=12\times 12$ to $L=36\times 36$ sites, which are illustrated in Fig. 3 for representative values of $U$ (U=1.0,0.7,0.6,5.0U=-1.0,-0.7,-0.6,5.0). One can see that in the case of $U\leq-1$ (the NO phase), we have found a broad maximum in the specific heat capacity curves which does not change with increasing $L$, and there is also no change in sublattice magnetizations, which equal zero. This can be attributed to the presence of a Schottky anomaly, which indicates a small number of discrete energy levels that dominate the system behavior and quantify the spacing between these levels [46]. For U>1U>-1 (where the F phase is found at low temperatures), we have found the typical $\lambda$-like singularity behavior of specific heat capacity, which increases its maximum value with the increasing $L$ as well as the rapid change in sublattice magnetizations from $|m|=0$ to $|m|=1$ with the decreasing temperature. To distinguish the order of these phase transitions, we have used the method by Challa et al. [47, 48, 49, 50], which is based on the energy distribution functions $P(E)$ for temperatures near phase transition. For $U$ smaller than $U\approx-0.65$ (temperatures smaller than $k_{B}T\approx 0.2$), the energy distribution function exhibits a two-peaked structure near the phase transition (as shown in Fig. 3 for U=0.7U=-0.7 and $k_{B}T\approx 0.19$), indicating a first-order phase transition from the NO phase to the F phase with decreasing temperature. This double-peak distribution indicates that states with two different energies of the system are equally probable at the transition temperature and one of them corresponds to the metastable solution in the vicinity of the transition line (however, the double-peak structure resembles a wide peak for model parameters shown). On the other hand, for higher $U$, only one peak in $P(E)$ is observed, which denotes a second-order phase transition (absence of metastability in the system) as shown for U=0.6U=-0.6 and $k_{B}T\approx 0.28$. We also presents the results for $U=5.0$, where we get the second-order transition at $k_{B}T\approx 0.604$, which is in very good accordance with the exact result for the Ising model ($k_{B}T\approx 0.607$) [10, 11]. However, the fluctuations are more visible due to higher temperatures and more Monte Carlo steps would provide smoother curves.
@@
Now we are ready to analyze the behavior of the system with changing $\mu$ (for $n\neq 1$). Exemplary $k_{B}T$$\mu$ phase diagrams for different values of $U$ are shown in Fig. 2. For all values of $U$ one sees the universal behavior of F–NO transition temperature, namely, the increase of $|\mu|$ leads to a decrease of the temperature. In the case of U=0.7U=-0.7, the transition remains of first-order across the entire range of the parameter $\mu$. Another behavior is observed for $U$ values larger than $U\approx-0.65$, where a second-order transition occurs near $\mu=0$ and, with increasing $|\mu|$, changes its order into a first-order transition (cf. diagrams for U=0.6U=-0.6 and U=0.5U=-0.5). With further increasing of onsite repulsion $U$, the region of the F phase occurrence extends and the first-order line becomes almost vertical (cf. $U=0.0$ case). It is remarkable that the temperature at which the F-NO transition changes its order (the tricritical point) is almost independent of $U$ and $\mu$, remaining around $k_{B}T\approx 0.18-0.20$, which is in an agreement with previous studies [32, 31].
In Fig. 4, we present dependence of several quantities with increasing temperature away half-filling to give an insight and justify the phase diagrams discussed above. The first column (U=0.7U=-0.7, $\mu=-0.15$) presents the dependencies in the NO phase with the broad peak in the specific heat capacity accompanied by zero sublattice magnetizations and continuous decrease in number of particles with decreasing temperature. The second and the third columns (U0.7U-0.7, $\mu=-0.10$ and U=0.6U=-0.6, $\mu=-0.18$) clearly indicate that, with increasing temperature, the first-order F-NO transition occurs, with a sharp (discontinuous) change of $m$ and $n$ at the transition temperature, and double peak-structure in the energy distribution $P(E)$ near the transition (which is more clearly visible than previously discussed example in Fig. 3). Finally, it is clearly seen that for U=0.6U=-0.6 and $\mu=-0.12$ (the fifth column) the F-NO transition is second-order. Note that the particle concentration $n$ approaches zero in the NO phase at the ground state, whereas in the F phase it goes to $1$ as it is expected for the exact ground-state results [32, 31].
It is worth to mention that the findings for J=1J=-1 are in a very good agreement, at least qualitatively, with the results obtained by the mean-field approach presented in [32, 31]. Obviously, the mean-field overestimates the transition temperatures. Moreover, the results for the ferromagnetic coupling are also very similar (qualitatively and quantitatively) to those obtained for the square lattice [36, 37, 38, 32]. The predicted transition temperatures for the triangular lattice ($z=6$) are slightly higher than those for the square lattice ($z=4$), which is due to the higher number of the nearest-neighbors in the former case.